# Fine-grained Attention in Hierarchical Transformers for Tabular Time-series

### v1

@@ 1. Introduction
As tabular data is bi-dimensional, and both rows and columns carry semantics, the transformer architectures available in literature for tabular time-series often present a hierarchical design. In a first stage, each row (or column) is encoded separately by aggregating the outputs of a transformer computing attention across its fields, as shown in Figure 1 – left and center. In the second stage, these encoded rows (or columns) are then passed to another transformer. The result of the second stage is the final encoding of the entire tabular time-series, which is typically processed by additional fully connected layers that specialize the model for solving a given downstream task, e.g., sequence classification. From a tabular perspective, this two-stage process first captures interactions between fields within a given dimension (row or column), and then interactions among those representations. While hierarchical architectures capture all tables’ dimensions, they don’t do that simultaneously, hence limiting visibility on more subtle cross-field relationships important for the downstream task, an effect which we empirically demonstrate in Appendix A. This suggests that a field-wise attention mechanism can be an appealing alternative to capture the intricate relationships between all the fields across the full tabular time-series, as illustrated in Figure 1 – right.
In this paper, we implement this mechanism, introducing $\mathsf{Fieldy}$, a novel architecture that combines row-wise and column-wise transformers in the first stage to learn field representations. These contextualized field representations are then merged, reshaped and passed to the second-stage transformer to produce the final encoding of the entire tabular time-series. Consequently, $\mathsf{Fieldy}$ enables fine-grained attention across all the fields composing a tabular time-series, while incorporating row-level and column-level information. We compare our solution against both state of the art transformer architectures and Machine Learning (ML) tree-based ensemble algorithms using two popular tabular datasets and show that $\mathsf{Fieldy}$ outperforms alternative methods.
@@ 2. Related work
Traditional tabular data modeling is row-based – each row is treated as a distinct input sample for which a prediction or a classification needs to be made. A variety of proposals address this type of data using both DL and ML methods. For instance, a recent study (Grinsztajn et al., 2022) explores DL approaches, including CNNs and transformers, as well as ML tree-based models like XGBoost and Random Forests, applied to tabular data. Although the study shows that gradient-boosted trees outperform deep learning algorithms on most datasets, FT-Transformer (Gorishniy et al., 2021) emerges as a promising architecture. In particular, FT-Transformer encodes each row by computing attention between its fields and then processes it, with a final fully connected layer. FT-Transformer relies on feature tokenizers to embed both categorical and numerical features that may appear in tabular data, which is an uncommon and challenging scenario for ML approaches. Note that this method is row-based but not hierarchical as each record is encoded separately. Alternatively, Tabbie (Iida et al., 2021) proposes to encode each table’s field by averaging representations of its row and its column. To do so, two transformers are tasked to encode each row and each column separately, in order to form “contextualized” fields representations by averaging their intersections. Once the full table has been encoded, each row is processed by a final fully connected layer. Note that Tabbie is not a hierarchical architecture, as it only operates at the field-level granularity, either contextualized by row or by column. While Tabbie is not adapted to tabular time-series, as it does not consider subsets of time-dependent rows, it is the closest to our proposal in the way it attends to row and column contexts.
Transformers for tabular time-series
In their most prevalent form, tabular time-series are a specific family of tabular data where records are interdependent and ordered or timestamped (e.g., with an explicit Timestamp field as in Table 1), often presenting an “entity” identifier used for grouping records into samples of interest. For example, Table 1 yields three distinct tabular time-series corresponding to three patients’ history, each of which could be the input to a machine learning model. In this case, a classic ML algorithm would simply take as input the concatenation of multiple records, hence returning to the traditional tabular data scenario. Conversely, more recent sequence models open the way for using hierarchical approaches to tabular time-series modeling. In this context, TabBERT (Padhi et al., 2021) is composed of a first-stage transformer that encodes each row in a tabular time-series, and a second-stage transformer that processes the encoded rows to generate a full sequence representation. This final representation is then passed to a fully connected layer, e.g., a classification head, to perform the downstream task. Given its design, we refer to this architecture as “row-based”. Alternatively, instead of encoding rows in the first stage, TabBERT may encode each column separately, and process these encoded columns in a second stage to generate the full tabular time-series representation. We refer to this inversion of TabBERT as “column-based”. Several works have extended this hierarchical approach to model tabular time-series, such as (Luetto et al., 2023) that extends it to heterogeneous tabular time-series i.e., with different number of columns, and (Zhang et al., 2023) that exploits time-deltas to better model time differences in the context of tabular time-series. Our work builds on hierarchical tabular time-series modeling, extending it with the ability to integrate field interactions across both rows and columns.
Our contribution
In this paper we set out to evaluate whether a finer cross-field attention, that simultaneously captures row-wise and column-wise interactions, provides more appropriate representations of tabular time-series, as opposed to a coarse aggregation of row or column embeddings. To properly isolate this effect and focus the comparison on this particular design choice, we compare this field-based attention approach to its row-based and column-based counterparts in similar conditions. In other terms, in this paper, we do not aim to introduce novel tokenizers or novel pre-training tasks which are orthogonal directions of research. Next, since our two-step approach relates any field to any other during the second stage, unlike row-based and column-based, we compare it also to a single-stage baseline where all the rows of interest are flattened and fed to the transformer, effectively resulting in a comparison point that links any field to any other. Finally, to complete the design space, another interesting comparison point we consider consists in constructing field embeddings that incorporate both rows and columns and then feeding the flattened sequences to a linear layer. This approach could be seen as a straightforward adaptation of Tabbie (Iida et al., 2021) to the context of tabular time series.
3. Methodology
In this section, we first dive into the details of current approaches and their limitations. We then present our field-wise attention mechanism that integrates both row-wise and column-wise interactions. We finally introduce the positional embeddings schemes used to encode table structure.
State of the art limitations
Tabular time-series modeling approaches considered in this paper are summarized in Table 2. Non-hierarchical baselines such as FT-Transformer and Tabbie are single-stage architectures designed for traditional (i.e., non sequential) tabular data. When adapting FT-Transformer to tabular time-series, the input sequence is flattened to form a single long row of repeated features at various time-steps, as depicted in Figure 2 (a). As this model computes attention across all the fields composing a single row, it captures relationships between all the fields across all rows and columns as a result of the input flattening. However, the downside of this flattening is that the FT-Transformer model is oblivious to the table structure. We consider this as a baseline transformer attending to all fields. Alternatively, when adapted to ingest tabular time-series, Tabbie limits its row-wise contextualization to a subset of the table: the rows composing the input sequence. As it relates each field to other fields present in the same row and column, Tabbie is able to capture relationships along both table axes. However, as depicted in Figure 2 (b), Tabbie is not equipped with a second-stage to relate all the fields outside of their original row and column.
Regarding hierarchical models, recall that two-stage architectures from the literature are either “row-based” or “column-based”. Thus, they condition the attention mechanism between distinct fields along a particular table axis in the first stage, as illustrated in Figure 2 (c). While these approaches enforce the tabular time-series row or column structure, they fail to capture relationships between fields across separate rows (or columns). For instance, considering Table 1, a row-based architecture cannot explicitly relate (by means of attention) a Disease field to a Therapy field if they belong to different rows, nor can it relate two Disease fields that belong to two different rows.
Field-wise attention
Two-stage approaches rely on the assumption that the Field transformer in the first stage is sufficiently powerful to extract expressive row/column representations for the Final transformer, rather than fostering the learning of fine-grained field relationships in the second stage. To capture relevant fields interations that may be beneficial to tabular time-series tasks, we introduce $\mathsf{Fieldy}$: a novel hierarchical transformer that combines both in-row and in-column attention, as well as cross-field attention.
As depicted in Figure 2 (d), we propose a two-stage architecture in which the first stage consists of two Field transformers that operate simultaneously: one is responsible to contextualize each field row-wise, and the other is responsible to contextualize each field column-wise. The resulting encoded rows and columns are then concatenated to constitute field representations. These contextualized cell representations are then passed through a fully connected layer to produce rich representations before being processed by a Final transformer, which would attend to all fields. Note that our first step resembles Tabbie; yet, it differs for two key design choices. First, while Tabbie creates a field embedding through a deterministic, simple average of the field’s row and column embeddings, $\mathsf{Fieldy}$ learns how to combine the two embeddings in the first stage. Second, unlike Tabbie which passes the embedded fields to the final fully connected layer (e.g., a classifier head), we adopt a second-stage transformer that relates all the field representations to each other. At last, the entire tabular time-series encoding coming from the Final transformer is finally processed by a fully connected layer, fine-tuned on a specific downstream task (e.g., sequence classification). In order to compare $\mathsf{Fieldy}$ against row-based and column-based architectures, we reduce the size of the Field transformers. We note that our approach requires increased computational effort, an aspect which we discuss in Section 5.
Positional encoding to capture table structure
Compared to row-based or column-based hierarchical approaches that model the table structure by design, $\mathsf{Fieldy}$ requires additional information. In the row-based architecture, the Field transformer implicitly establishes the table’s horizontal structure by computing attention horizontally across all the fields of the same row. The Final transformer then enforces the table vertical structure as it ingests a sequence of per-row representations. Conversely, in the column-based architecture, the Field transformer is provided with the table’s vertical structure, as its input is a sequence of fields from distinct rows. Then, the Final transformer enforces the table horizontal structure by attending between per-column representations. For these two approaches, it is only the position of the rows (or columns) that is unknown from the model, but their delimitation is apparent by design.
On the other hand, the design of our proposed two-stage architecture does not incorporate the table structure by default, as the Final transformer ingests a long sequence composed of all the contextualized fields. Thus, without additional information, the Final transformer can only access a bag of fields, being oblivious to the original rows or columns they come from. Therefore, we incorporate row and column positional embeddings (Devlin et al., 2018). Before being passed to the Final transformer, each contextualized field is augmented (by means of element-wise addition) with two embeddings: one carrying its original row position and one carrying its original column index. To ensure a fair comparison between architectures, we also add these row position and column index embeddings to all the other models. We discuss the effect of these additions in Section 5.
@@ 4. Evaluation
In this section, we first detail the datasets and models configurations considered to compare transformers on tabular time-series. Then, we present the results of our evaluation and expand our analysis with an ablation study.
@@ 4.1. Datasets
The UCI Beijing Pollution Dataset (Chen, 2019) consists in predicting air pollution particles from 12 sites located in Beijing. This dataset has been used to evaluate row-based hierarchical transformers in (Padhi et al., 2021; Luetto et al., 2023). It is a multi-regression task taking as input 11 features (such as temperature, pressure, etc.) measured on an hourly basis during 10 time-steps. The labels to predict correspond to the PM2.5 and PM10 concentrations for each time-step, i.e., 20 labels for each input sequence. We replicate the pre-processing described in (Padhi et al., 2021) i.e., we discretize numerical variables using 50 quantiles and normalize the targets. After data exploration, we choose to include 5 additional features not present in the pre-processing from related work. These engineered features correspond to the measurement site name, the hour of the measurement, the day of the month, the weekday, the month, and the year. Finally, we remove 4% of outliers when PM10 > PM2.5 as mentioned in (Wu et al., 2018). As in prior work, the evaluation metric is defined as the RMSE averaged across the concentration targets. In total, this dataset contains 67K tabular time-series. The pre-training dataset consists in the same dataset excluding the labels (more details on data splits in Appendix C).
@@
The PKDD’99 Bank transactions dataset (Berka, 1999) contains real transaction records for 4,500 clients of a Czech bank. It has been used in (Luetto et al., 2023) to evaluate hierarchical transformers. The considered task consists in predicting if a client will default its loan based on his prior transactions. Six input features describe each transaction (amount, type, etc.) and the label is binary, i.e., one for each clients’ input sequence. As for the pollution dataset, we pre-process the data similarly to prior work (Luetto et al., 2023) i.e., using 50 quantiles to discretize each numerical feature and splitting the timestamp into three fields: day, month and year. Note that we include an additional weekday feature that proved meaningful during data exploration. As in (Luetto et al., 2023), we segregate the data into 3,818 clients with unlabeled transactions for pre-training, and 682 clients with labeled transactions for fine-tuning. In order to increase the dataset size, we consider any sequence of 10 consecutive transactions for each client, while (Luetto et al., 2023) considered only the last 150 client’s transactions. We thus obtain 5K tabular time-series for fine-tuning, instead of only 682. The evaluation metric is defined as the Average Precision score, to take into account the class imbalance of this dataset (more details on data splits in Appendix C).

### v2

@@ Abstract.
journalyear: 2024copyright: noneconference: 10th Mining and Learning from Time Series Workshop (MiLeTS); August 26, 2024; Barcelona, Spainbooktitle: ACM SIGKDD’24 – 10th Mining and Learning from Time Series Workshop (MiLeTS), August 26, 2024, Barcelona, Spain
@@ 1. Introduction
As tabular data is bi-dimensional, and both rows and columns carry semantics, the transformer architectures available in literature for tabular time-series often present a hierarchical design. In a first stage, each row (or column) is encoded separately by aggregating the outputs of a transformer computing attention across its fields, as shown in Figure 1 – left and center. In the second stage, these encoded rows (or columns) are then passed to another transformer. The result of the second stage is the final encoding of the entire tabular time-series, which is typically processed by additional fully connected layers that specialize the model for solving a given downstream task, e.g., sequence classification. From a tabular perspective, this two-stage process first captures interactions between fields within a given dimension (row or column), and then interactions among those representations.
While hierarchical architectures capture all table dimensions, they don’t do that simultaneously, hence limiting visibility on more subtle cross-field relationships important for the downstream task. In Appendix A, we empirically demonstrate that this shortcoming hinders learning interactions between fields across separate rows, due to the coarse-grained aggregation performed in the second stage. This suggests that a field-wise attention mechanism can be an appealing alternative to capture the intricate relationships between all the fields across the full tabular time-series, as illustrated in Figure 1 – right. In this paper, we implement this mechanism, introducing $\mathsf{Fieldy}$, a novel architecture that combines row-wise and column-wise transformers in the first stage to learn field representations. These contextualized field representations are then merged, reshaped and passed to the second-stage transformer to produce the final encoding of the entire tabular time-series. Consequently, $\mathsf{Fieldy}$ enables fine-grained attention across all the fields composing a tabular time-series, while incorporating row-level and column-level information. We compare our solution against both state of the art transformer architectures and Machine Learning (ML) tree-based ensemble algorithms using two popular tabular datasets and show that $\mathsf{Fieldy}$ outperforms alternative methods.
@@ 2. Related work
In traditional tabular data modeling, each row is treated as a distinct input sample for which a prediction or a classification needs to be made. A variety of proposals address this type of data using both DL and ML methods. For instance, a recent study (Grinsztajn et al., 2022) explores DL approaches, including CNNs and transformers, as well as ML tree-based models like XGBoost and Random Forests, applied to tabular data. Although the study shows that gradient-boosted trees outperform deep learning algorithms on most datasets, FT-Transformer (Gorishniy et al., 2021) emerges as a promising architecture. In particular, FT-Transformer encodes each row by computing attention between its fields and then processes it, with a final fully connected layer. FT-Transformer relies on feature tokenizers to embed both categorical and numerical features that may appear in tabular data, which is an uncommon and challenging scenario for ML approaches. Note that this method is row-based but not hierarchical as each record is encoded separately. Alternatively, Tabbie (Iida et al., 2021) proposes to encode each table field by averaging representations of its row and its column. To do so, two transformers are tasked to encode each row and each column separately, in order to form “contextualized” fields representations by averaging their intersections. Once the full table has been encoded, each row is processed by a final fully connected layer. Note that Tabbie is not a hierarchical architecture, as it only operates at the field-level granularity, either contextualized by row or by column. While Tabbie is not adapted to tabular time-series, as it does not consider subsets of time-dependent rows, it is the closest to our proposal in the way it attends to row and column contexts.
@@
Transformers for tabular time-series
In their most prevalent form, tabular time-series are a specific family of tabular data where records are interdependent and ordered or timestamped (e.g., with an explicit Timestamp field as in Table 1), often presenting an “entity” identifier used for grouping records into samples of interest. For example, Table 1 yields three distinct tabular time-series corresponding to three patients’ history, each of which could be the input to a machine learning model. In this case, a classic ML algorithm would simply take as input the concatenation of multiple records, hence returning to the traditional tabular data scenario. Conversely, more recent sequence models open the way for using hierarchical approaches to tabular time-series modeling. In this context, TabBERT (Padhi et al., 2021) is composed of a first-stage transformer that encodes each row in a tabular time-series, and a second-stage transformer that processes the encoded rows to generate a full sequence representation. This final representation is then passed to a fully connected layer, e.g., a classification head, to perform the downstream task. Given its design, we refer to this architecture as “row-based”. Alternatively, instead of encoding rows in the first stage, a variation of TabBERT may encode each column separately, and process these encoded columns in a second stage to generate the full tabular time-series representation. We refer to this inversion of TabBERT as “column-based”. Several works have extended this hierarchical approach to model tabular time-series, such as (Luetto et al., 2023) that extends it to heterogeneous tabular time-series, i.e., with different number of columns, and (Zhang et al., 2023) that exploits time-deltas to better model time differences in the context of tabular time-series. Our work builds on hierarchical tabular time-series modeling, extending it with the ability to integrate field interactions across both rows and columns simultaneously.
Our contribution
In this paper we set out to evaluate whether a hierarchical model with a finer cross-field attention provides more appropriate representations of tabular time-series, as opposed to a coarse aggregation of row or column embeddings. In a nutshell, we propose an architecture that simultaneously captures row-wise and column-wise interactions in a first stage, to learn contextualized field representations that are related to one another in a second stage. To properly isolate this effect and focus the comparison on this particular design choice, we compare this field-based attention approach to its row-based and column-based counterparts in similar conditions. In other terms, in this paper, we do not aim to introduce novel tokenizers or novel pre-training tasks which are orthogonal directions of research. Additionally, since our approach relates any field to any other during the second stage, unlike row-based and column-based, we compare it also to a single-stage baseline where all the rows of interest are flattened and fed to a unique transformer. This approach can be seen as an adaptation of FT-Transformer for tabular time-series, effectively resulting in a comparison point that links any field to any other. Finally, to complete the design space, another interesting comparison point we consider consists in constructing contextualized field embeddings that incorporate both rows and columns and then feeding the flattened sequence to a linear layer. This approach could be seen as a straightforward adaptation of Tabbie to the context of tabular time series.
3. Methodology
In this section, we first dive into the details of current approaches and their limitations. We then present our field-based hierarchical approach that integrates both row-wise and column-wise interactions. We finally introduce the positional embeddings schemes used to encode table structure.
State of the art limitations
Tabular time-series modeling approaches considered in this paper are summarized in Table 2. Non-hierarchical baselines such as FT-Transformer and Tabbie are single-stage architectures designed for traditional (i.e., non sequential) tabular data. When adapting FT-Transformer to tabular time-series, the input sequence is flattened to form a single long row of repeated features at various time-steps, as depicted in Figure 2 (a). As this model computes attention across all the fields composing a single row, it captures relationships between all the fields across all rows and columns as a result of the input flattening. However, the downside of this flattening is that this adaptation of the FT-Transformer model is oblivious to the table structure. We consider this as a baseline transformer attending to all fields. Alternatively, when adapting Tabbie to ingest tabular time-series, we limit its row-wise contextualization to a subset of the table: the rows composing the input sequence. As it relates each field to other fields present in the same row and column, Tabbie is able to capture relationships along both table axes. However, as depicted in Figure 2 (b), this approach is not equipped with a second-stage to relate all the fields outside of their original row and column.
Regarding hierarchical models, recall that two-stage architectures from the literature are either “row-based” or “column-based”. Thus, they condition the attention mechanism between distinct fields along a particular table axis in the first stage, as illustrated in Figure 2 (c). While these approaches enforce the tabular time-series row (or column) structure, they fail to capture relationships between fields across separate rows (or columns). For instance, considering Table 1, a row-based architecture cannot explicitly relate, i.e., by means of attention, a Disease field to a Therapy field if they belong to different rows, nor can it directly relate two Disease fields that belong to two different rows. As shown empirically in Appendix A, this limits the ability of hierarchical models from the literature to learn fine-grained relationships at the field level that might be relevant for a given downstream task.
Field-wise attention
Two-stage approaches rely on the assumption that the Field transformer in the first stage is sufficiently powerful to extract expressive row/column representations for the Final transformer, rather than fostering the learning of fine-grained field relationships in the second stage. To capture relevant fields interations that may be beneficial to tabular time-series tasks, we introduce $\mathsf{Fieldy}$: a novel hierarchical transformer that combines both in-row and in-column attention, as well as cross-field attention. As depicted in Figure 2 (d), we propose a two-stage architecture in which the first stage consists of two Field transformers that operate simultaneously: one is responsible to contextualize each field row-wise, and the other is responsible to contextualize each field column-wise. The resulting encoded rows and columns are then concatenated to constitute field representations. These contextualized cell representations are then passed through a fully connected layer to produce rich representations before being processed by a Final transformer, which would attend to all fields. Note that our first stage resembles Tabbie; yet, it differs for two key design choices. First, while Tabbie creates a field embedding through a deterministic simple average of the field’s row and column embeddings, $\mathsf{Fieldy}$ learns how to combine the two embeddings in the first stage. Second, unlike Tabbie which passes the embedded fields to the final fully connected layer (e.g., a classifier head), we adopt a second-stage transformer that relates all the field representations to each other. At last, the entire encoded tabular time-series coming from the Final transformer is processed by a fully connected layer, fine-tuned on a specific downstream task (e.g., sequence classification). In order to fairly compare $\mathsf{Fieldy}$ against row-based and column-based architectures, we reduce the size of its Field transformers to reach the same model size. Nonetheless, we note that $\mathsf{Fieldy}$ requires increased computational effort, an aspect which we discuss in Section 5.
Positional encoding to capture table structure
Compared to row-based or column-based hierarchical approaches that model the table structure by design, $\mathsf{Fieldy}$ requires additional information. In the row-based architecture, the Field transformer implicitly establishes the table’s horizontal structure by computing attention horizontally across all the fields of the same row. The Final transformer then enforces the table vertical structure as it ingests a sequence of per-row representations. Conversely, in the column-based architecture, the Field transformer is provided with the table’s vertical structure, as its input is a sequence of fields from distinct rows. Then, the Final transformer enforces the table horizontal structure by attending between per-column representations. For these two approaches, it is only the position of the rows (or columns) that is unknown from the model, but their delimitation is apparent by design. On the other hand, the design of $\mathsf{Fieldy}$ does not incorporate the table structure by default, as the Final transformer ingests a long sequence composed of all the contextualized fields. Thus, without additional information, the Final transformer can only access a bag of fields, being oblivious to the original rows or columns they come from. Therefore, we incorporate row and column positional embeddings (Devlin et al., 2018). Before being passed to the Final transformer, each contextualized field is augmented (by means of element-wise addition) with two embeddings: one carrying its original row position and one carrying its original column index. To ensure a fair comparison between architectures, we also add these row position and column index embeddings to all the other models. We discuss the effect of these additions in Section 4.
@@ 4. Evaluation
In this section, we first detail the datasets and models configurations considered to compare transformers architectures on tabular time-series. Then, we present the results of our evaluation and expand our analysis with an ablation study.
@@ 4.1. Datasets
The UCI Beijing Pollution Dataset (Chen, 2019) consists in predicting air pollution particles from 12 sites located in Beijing. This dataset has been used to evaluate row-based hierarchical transformers in (Padhi et al., 2021; Luetto et al., 2023). It is a multi-regression task taking as input 10 features (such as temperature, pressure, etc.) measured on an hourly basis during 10 time-steps. The labels to predict correspond to the PM2.5 and PM10 concentrations for each time-step, i.e., 20 labels for each input sequence. We replicate the pre-processing described in (Padhi et al., 2021), i.e., we discretize numerical variables using 50 quantiles and normalize the targets. After data exploration, we choose to include 6 additional features not present in the pre-processing from related work. These engineered features correspond to the measurement site name, the hour of the measurement, the day of the month, the weekday, the month, and the year. Finally, we remove 4% of outliers when PM10 > PM2.5 as mentioned in (Wu et al., 2018). As in prior work, the evaluation metric is defined as the RMSE averaged across the concentration targets. In total, this dataset contains 67K tabular time-series. The pre-training dataset consists of the same dataset excluding the labels (more details on data splits in Appendix C).
@@
The PKDD’99 Bank transactions dataset (Berka, 1999) contains real transaction records for 4,500 clients of a Czech bank. It has been used in (Luetto et al., 2023) to evaluate hierarchical transformers. The considered task consists in predicting if a client will default its loan based on his prior transactions. Six input features describe each transaction (amount, type, etc.) and the label is binary, i.e., one for each clients’ input sequence. As for the pollution dataset, we pre-process the data similarly to prior work (Luetto et al., 2023), i.e., using 50 quantiles to discretize each numerical feature and splitting the timestamp into three fields: day, month and year. Note that we include an additional weekday feature that proved meaningful during data exploration. As in (Luetto et al., 2023), we segregate the data into 3,818 clients with unlabeled transactions for pre-training, and 682 clients with labeled transactions for fine-tuning. In order to increase the dataset size, we consider any sequence of 10 consecutive transactions for each client, while (Luetto et al., 2023) considered only his last 150 transactions. We thus obtain 5K tabular time-series for fine-tuning, instead of only 682. The evaluation metric is defined as the Average Precision (AP) score to take into account the class imbalance of this dataset (more details on data splits in Appendix C).